OpenStudy (anonymous):

HELP! CD is the perpendicular bisector of AD. Find AC AD and AE ill have pic below

OpenStudy (anonymous):



OpenStudy (anonymous):

Pythagorean theorem: a^2 + b^2 = c^2 where a, b are regular sides and c is the hypotenuse. You have all the info you need.

OpenStudy (anonymous):

alright alright tahtnk you

OpenStudy (anonymous):

@saifoo.khan think you could help me out :)

OpenStudy (anonymous):

How far did you get? Show me your work and we can do this together.

OpenStudy (anonymous):

im kinda confused about that with all the x's and y's so do I have to solve for them first befor I can do anything?

OpenStudy (anonymous):

You do not have to solve for x and y. You can figure out the problem as a system of equations.

OpenStudy (anonymous):

Ok so the do I take9y-6^2+3x=c^2

OpenStudy (anonymous):

to find AC and AD

OpenStudy (anonymous):

(9y-6)^2 + (3x)^2 = c^2. 81y2 - 108y + 36 + 9x^2 = c^2 Then you basically take the square root of that, which I can't factor. Is there a way to find x or y? Do you have a numerical reference?

OpenStudy (anonymous):

I see you have 5x = 6y up there, is that a typo?

OpenStudy (anonymous):

yes it supposta be 5x+6y

OpenStudy (anonymous):

so the answers I get for lthe problems you did will be AC AD?

OpenStudy (anonymous):

well it depends what format you need them in. if your teacher wants polynomials give him polynomials.

OpenStudy (anonymous):

so I got 81y-36+9=c^2 and 6561y-11664y+36+81......I don't have a teacher im homeschooled and this is all the book gives me to solve